# 3rd Place Report of LSVOS 2025 MeViS Track: Sa2VA-i: Improving Sa2VA Results with Consistent Training and Inference

### v1

@@ Abstract
Sa2VA is a recent model for language-guided dense grounding in images and video that achieves state-of-the-art results on multiple segmentation benchmarks and that has become widely popular. However, we found that Sa2VA does not perform according to its full potential for referring video object segmentation tasks. We identify inconsistencies between training and inference procedures as the key factor holding it back. To mitigate this issue, we propose an improved version of Sa2VA, Sa2VA-i, that rectifies these issues and improves the results. In fact, Sa2VA-i sets a new state of the art for multiple video benchmarks and achieves improvements of up to +11.6 $\mathcal{J}$$\mathcal{F}$ on MeViS, +1.4 on Ref-YT-VOS, +3.3 on Ref-DAVIS and +4.1 on ReVOS using the same Sa2VA checkpoints. With our fixes, the Sa2VA-i-1B model even performs on par with the original Sa2VA-26B model on the MeViS benchmark. We hope that this work will show the importance of seemingly trivial implementation details and that it will provide valuable insights for the referring video segmentation field. We provide the code and updated models at github.com/kumuji/sa2va-i.
@@ 1 Introduction
Accurately localizing and tracking objects in videos is an important area in computer vision with multiple real-world applications, ranging from video editing [5] and robotics [28] to medical [20] and biological [51] applications. Over the years, a large volume of work [47, 50, 32, 29, 2, 31] proposed datasets, benchmarks, and model architectures that focus on segmenting and tracking objects. The recent surge in vision-language models [27, 1] has transformed the landscape of video segmentation research, particularly in segmenting and tracking objects based on complex text descriptions. This task, called Referring (and later extended to Reasoning) Video Object Segmentation (RVOS), has recently attracted significant attention, with multiple new datasets [11, 6, 48, 3] and methods [4, 48, 52] emerging.
Earlier architectures that addressed RVOS [7, 44] relied on vision-language models such as CLIP [34] to extract features from the image frames and text prompts, coupled with a segmentation model that predicted masks. While effective for short text prompts, this approach struggles with detailed text descriptions requiring logical reasoning. Therefore, recent works [46, 4, 23, 48, 30] employ Multimodal Large Language Models (MLLM), which are given video features and a text prompt as input and provide segmentation masks as an output. The prevailing practice is to train the MLLM to predict a special [SEG] token that represents the object to be segmented, and to use a segmentation model such as SAM or SAM2 [22, 36] to predict a segmentation mask based on this token. The current state-of-the-art model, Sa2VA [52], adopts this paradigm, but scales the approach with more data and uses more recent MLLMs.
Sa2VA introduces a prompting strategy that uses the streaming memory of SAM2 to segment objects in a video. Despite achieving great results on multiple benchmarks, we identify a critical discrepancy between the training and inference procedure. During training, only the SAM2 mask decoder is finetuned, while the memory encoder and attention components remain frozen and unused. This is inconsistent with the inference procedure, because these frozen components are used during inference, operating on feature representations from a finetuned mask decoder, which they were never optimized to handle. Thus, although the memory prompting mechanism is theoretically sound, its performance is suboptimal due to this mismatch between training and inference, particularly on challenging video segmentation benchmarks such as MeViS and ReVOS [11, 48].
@@ 2 Related Work
Video Segmentation with MLLMs  With recent advances in vision-language research, multiple referring image segmentation methods have been proposed [46, 23, 54, 35], sparking interest in segmenting and tracking objects in videos, i.e., Reasoning and Referring Video Object Segmentation (RVOS). Although initial architectures for RVOS involved using text and vision encoders coupled with a Transformer decoder to regress masks [49, 7, 44, 6, 39], recent approaches [4, 48, 52] adopt MLLM-based architectures to better reason about complex text prompts and to combine multiple tasks. This is usually done by extending a Multimodal LLM architecture with a downstream segmentation model and by predicting a special segmentation token for each target object. This token is then input into the segmentation model together with video frame features to predict masks. For the video segmentation task, the image-first PSALM model [54] can be conditioned on a previous mask to segment the next frame. Similarly, TrackGPT [58] processes a single frame at a time. VISA [48] predicts a single token for a carefully selected keyframe in a video and tracks that corresponding object. VideoLISA [4] uses SAM [22] and reuses a single token for multiple frames. More recently, VideoGLaMM [30] and Sa2VA [52] use the SAM2 [36] segmentation model, but do not finetune the memory components. In this work, we identify that Sa2VA operates differently during inference and training, which harms the performance. To solve this, we present Sa2VA-i, which ensures consistent training and inference and considerably outperforms Sa2VA.
@@
Works that use Sa2VA  Sa2VA has rapidly become a foundational model for language-grounded segmentation, and has been widely adopted and extended across diverse research efforts. For instance, its capabilities are leveraged for dataset creation, with AnimeShooter [33] using Sa2VA to extract masks of animated characters. Furthermore, several new benchmarks, including EASG-Bench [37], FinChart-Bench [41] and V-STaR [9], evaluate Sa2VA as a state-of-the-art method for spatial grounding. The model is also a baseline for comparison to SpatialReasoner-R1 [40], Omni-R1 [56], SeC [55], and DenseWorld-1M [24], EgoMask [25]. InterRVOS [18], Dense360 [57], UniBiomed [45], DeSa2VA [19], and SAMA [42] are built directly on top of Sa2VA. The practical utility of Sa2VA is evident in applications such as Gondola [8] and FineGrasp [15], which use it to segment objects for grasping. Furthermore, Sa2VA prominence was highlighted during the PVUW MeViS Challenge at CVPR 2025 [13], where the first [16] and third [53] place solutions were based on Sa2VA. The frequent use of Sa2VA in the literature highlights the relevance of this work in which we identify and solve the inference-training inconsistency in Sa2VA. With Sa2VA-i, we provide a stronger base model for various benchmarks, applications and downstream tasks. We encourage authors to apply our suggested improvements to ensure proper comparisons and improve their results.
@@ 3.1 Preliminaries
SAM and SAM2  SAM [22] is an interactive image segmentation model that consists of image encoder, prompt encoder and mask decoder. Box, point or mask inputs from the user are encoded by the prompt encoder into a single feature vector. Using this feature vector, as well as image features encoded by the image encoder, the mask decoder generates a segmentation output for the selected object. For interactive video segmentation, SAM2 [36] augments the original SAM architecture with several memory components. SAM2 processes videos in a streaming fashion, i.e., frame by frame. Like SAM, SAM2 segments objects based on a box prompt, mask prompt or point prompt, which are provided for one of the video frames in this case. Given an input frame, SAM2 first generates per-frame features with its image encoder. Then, SAM2 applies memory attention to condition these per-frame features on the features from earlier frames in which an object was segmented, which were stored in a memory bank. Subsequently, the mask decoder predicts a segmentation mask based on the conditioned frame and the input prompts. Finally, the image and object features are encoded using a memory encoder and stored in the memory bank such that they can be used to condition future frames.
@@
For referring video segmentation, Sa2VA also predicts only a single [SEG] token, and reuses this token to predict masks for multiple video frames, similar to VideoLISA [4]. Due to the limited context window of the MLLM, Sa2VA samples $T$ video frames, only feeds the visual tokens for those $T$ frames through the MLLM, and makes segmentation predictions with SAM2 only for those frames. To obtain predictions for all $I>T$ frames in a video, Sa2VA applies post-processing using the same SAM2 decoder.
@@
Method MLLM Frames MeViS [11] Ref-YT-VOS [39] Ref-DAVIS17 [21]
@@
LISA [23] LLaVA-7B 37.2 35.1 39.4 53.9 53.4 54.3 64.8 62.2 67.3
LISA [23] LLaVA-13B 37.9 35.8 40.0 54.4 54.0 54.8 66.0 63.2 68.8
TrackGPT [58] LLaVA-7B 40.1 37.6 42.6 56.4 55.3 57.4 63.2 59.4 67.0
TrackGPT [58] LLaVA-13B 41.2 39.2 43.1 59.5 58.1 60.8 66.5 62.7 70.4
PSALM [54] Phi-1.5 68.8 65.9 71.7
VISA [48] Chat-UniVi-7B 43.5 40.7 46.3 61.5 59.8 63.2 69.4 66.3 72.5
VISA [48] Chat-UniVi-13B 44.5 41.8 47.1 63.0 61.4 64.7 70.4 67.0 73.8
VideoLISA [4] LLaVa-Phi-3-V-3.8B 44.4 41.3 47.6 63.7 61.7 65.7 68.8 64.9 72.7
InstructSeg [43] Mipha-3B 67.5 65.4 69.5 71.1 67.3 74.9
VideoGLaMM [30] Phi3-Mini-3.8B 45.2 42.0 48.2
SAMWISE [10] RoBERTa 49.5 46.6 52.4 69.2 67.8 70.6 70.6 67.4 74.5
VRS-HQ [17] Chat-UniVi-13B 50.9 48.0 53.7 71.0 69.0 73.1 74.4 71.0 77.9
GLUS [26] LLaVA-7B 51.3 48.5 54.2 67.3 65.5 69.0
MPG-SAM 2 [38] BEiT 53.7 50.7 56.7 73.9 71.7 76.1 72.4 68.8 76.0
Sa2VA [52] InternVL2.5-1B 5 47.0 68.0 69.5
Sa2VA [52] InternVL2.5-4B 5 46.4 43.3 49.5 71.3 69.1 73.5 73.7 69.6 77.8
Sa2VA [52] InternVL2.5-8B 5 51.5 72.3 75.9
Sa2VA [52] InternVL2.5-26B 5 52.1 75.1 78.6
@@
Table 2: Reasoning Video Segmentation on ReVOS [48]. Sa2VA-i outperforms Sa2VA and other methods by a large margin, obtaining state-of-the-art performance.
@@
VISA [48] Chat-UniVi-7B 39.2 36.7 41.7 52.9 51.1 54.7 46.1 43.9 48.2
VISA [48] Chat-UniVi-13B 40.9 38.3 43.5 54.1 52.3 55.8 47.5 45.3 49.7
InstructSeg [43] Mipha-3B 51.9 49.2 54.7 57.0 54.8 59.2 54.5 52.0 56.9
VRS-HQ [17] Chat-UniVi-13B 56.8 54.1 59.4 63.3 61.1 65.5 60.0 57.6 62.5
Sa2VA [52] InternVL2.5-4B 56.2 53.2 59.1 63.0 60.5 65.5 59.6 56.8 62.3
@@ 4.1 Experimental Setup
Evaluation  We evaluate Sa2VA-i using multiple popular benchmarks for video-level referring and reasoning segmentation [48, 21, 11, 39]. We use the standard $\mathcal{J}\&\mathcal{F}$ metric, which is the geometric mean of the Jaccard Score, $\mathcal{J}$, and the Boundary F-Score, $\mathcal{F}$ [31]. We use the MeViS dataset [11] for ablations, and report scores obtained from the public evaluation server on the val split.
@@ 4.2 Comparison to State of the Art
We compare Sa2VA-i with existing state-of-the-art referring video segmentation methods on multiple benchmarks. As shown in Tab. 1 and Tab. 2, Sa2VA-i achieves state-of-the-art results on all reported benchmarks and improves upon Sa2VA results using the same checkpoints and the same number of frames. Sa2VA-i’s performance improvement is particularly substantial on the challenging MeVIS benchmark [11]. It scores +10.2 $\mathcal{J}$$\mathcal{F}$ higher for the 4B model, +8.0 $\mathcal{J}$$\mathcal{F}$ for the 8B model, and +11.1 $\mathcal{J}$$\mathcal{F}$ for the 26B model. We attribute performance improvements to the improved consistency between training and inference, as well as better frame sampling. Additionally, we observe that the number of sampled frames during testing has a large influence on the results, but that the optimal value depends on the dataset.
@@ 4.3 Experiments
Table 3: Consistent Training and Inference. Using Sa2VA’s original model weights, we show gradual improvements that come from using our improved Sa2VA-i version with consistent training and inference. We further improve performance by using uniform sampling during inference instead of sampling first frames. Evaluation on MeViS val [11].
@@
Table 4: Frame Sampling. We evaluate the impact of different frame sampling strategies for both Sa2VA and Sa2VA-i. Uniform sampling consistently improves performance, with the largest gains observed when training and inference sampling are matched. Evaluation on MeViS val [11].
@@
Frame Sampling  We experiment with different frame sampling strategies during both training and inference, and report the results in Tab. 4. Similar to other works [16, 53], we observe improved performance when using uniform sampling during inference. To ensure training-inference consistency and to see the possible gains from training the model with uniform sampling, we train a 4B-parameter model with 8 frames sampled randomly and uniformly. During training with uniform sampling, we apply a random offset from the start to increase variation in the frames seen during training. We observe that uniform sampling and consistent sampling strategy yields even better performance, even for the original Sa2VA model.
Table 5: Number of Sampled Frames. Using Sa2VA’s original model weights, we show that gradual improvements can be obtained by using more frames during inference. Performance is progressively improved, with diminishing returns when using more then 12 frames. Evaluation on MeViS val [11].
@@
Number of Sampled Frames  Tab. 5 evaluates the impact of increasing the number of sampled frames that are fed to the MLLM and used to make the initial segmentation predictions, with the same model that has seen 5 frames during training. We find that increasing the number of sampled frames improves performance by non-negligible margins. In particular, the MeVIS performance can be improved by +1.2 $\mathcal{J}$$\mathcal{F}$ when using 12 sampled frames instead of 5. These findings are similar to the results reported in [16], where Sa2VA obtains higher scores when using more frames. However, while the original Sa2VA appears to benefit from using more frames without performance degradation, we observe diminishing returns for Sa2VA-i after sampling more than 12 frames.
@@
Table 6: Overprompting on Ref-DAVIS17 [21]. We vary the number of frames processed by the MLLM and the number of initially predicted masks used to prompt SAM2 for mask propagation. On this dataset, where the target is typically visible in the first frame, performance improves as the MLLM observes more frames, while prompting SAM2 with multiple initial masks degrades performance.
@@
The 7th LSVOS challenge at ICCV 2025 features two distinct tracks: a Video Object Segmentation (VOS) track evaluated on the complex LVOS and MOSE datasets [12, 14], and a Referring Video Object Segmentation (RVOS) track utilizing the motion-focused MeViS benchmark [11]. As shown in Tab. 7, our method achieved third place in the competition on the RVOS track, demonstrating strong performance in addressing these challenging scenarios. We achieve this result by using 12 frames during inference.

### v2

# Sa2VA-i: Improving Sa2VA Results with Consistent Training and Inference

@@ Abstract
Sa2VA is a recent model for language-guided dense grounding in images and video that achieves state-of-the-art results on multiple segmentation benchmarks and that has become widely popular. However, we found that Sa2VA does not perform according to its full potential for referring video object segmentation tasks. We identify inconsistencies between training and inference procedures as the key factor holding it back. To mitigate this issue, we propose an improved version of Sa2VA, Sa2VA-i, that rectifies these issues and improves the results. In fact, Sa2VA-i sets a new state of the art for multiple video benchmarks and achieves improvements of up to +11.6 $\mathcal{J}$$\mathcal{F}$ on MeViS, +1.4 on Ref-YT-VOS, +3.3 on Ref-DAVIS and +4.1 on ReVOS using the same Sa2VA checkpoints. With our fixes, the Sa2VA-i-1B model even performs on par with the original Sa2VA-26B model on the MeViS benchmark. We hope that this work will show the importance of seemingly trivial implementation details and that it will provide valuable insights for the referring video segmentation field. We provide the code and updated models at https://github.com/kumuji/sa2va-i.
@@ 1 Introduction
Accurately localizing and tracking objects in videos is an important area in computer vision with multiple real-world applications, ranging from video editing [bekuzarov2023xmemplus] and robotics [lu2023selfsupervised] to medical [khader2024medical] and biological [yang2024deep] applications. Over the years, a large volume of work [xu2018youtubevos, yang2019youtubevis, qi2022ovis, miao2022vipseg, athar2023burst, perazzi2016davis] proposed datasets, benchmarks, and model architectures that focus on segmenting and tracking objects. The recent surge in vision-language models [liu2023llava, alayrac2022flamingo] has transformed the landscape of video segmentation research, particularly in segmenting and tracking objects based on complex text descriptions. This task, called Referring (and later extended to Reasoning) Video Object Segmentation (RVOS), has recently attracted significant attention, with multiple new datasets [ding2023mevis, bellver2020refvos, yan2024visa, athar2025vicas] and methods [bai2024videolisa, yan2024visa, yuan2025sa2va] emerging.
Earlier architectures that addressed RVOS [botach2022mttr, wu2022referformer] relied on vision-language models such as CLIP [radford2021clip] to extract features from the image frames and text prompts, coupled with a segmentation model that predicted masks. While effective for short text prompts, this approach struggles with detailed text descriptions requiring logical reasoning. Therefore, recent works [xia2024gsva, bai2024videolisa, lai2023lisa, yan2024visa, munasinghe2024videoglamma] employ Multimodal Large Language Models (MLLM), which are given video features and a text prompt as input and provide segmentation masks as an output. The prevailing practice is to train the MLLM to predict a special [SEG] token that represents the object to be segmented, and to use a segmentation model such as SAM or SAM2 [kirillov2023sam, ravi2025sam2] to predict a segmentation mask based on this token. The current state-of-the-art model, Sa2VA [yuan2025sa2va], adopts this paradigm, but scales the approach with more data and uses more recent MLLMs.
Sa2VA introduces a prompting strategy that uses the streaming memory of SAM2 to segment objects in a video. Despite achieving great results on multiple benchmarks, we identify a critical discrepancy between the training and inference procedure. During training, only the SAM2 mask decoder is finetuned, while the memory encoder and attention components remain frozen and unused. This is inconsistent with the inference procedure, because these frozen components are used during inference, operating on feature representations from a finetuned mask decoder, which they were never optimized to handle. Thus, although the memory prompting mechanism is theoretically sound, its performance is suboptimal due to this mismatch between training and inference, particularly on challenging video segmentation benchmarks such as MeViS and ReVOS [ding2023mevis, yan2024visa].
@@ 2 Related Work
Video Segmentation with MLLMs  With recent advances in vision-language research, multiple referring image segmentation methods have been proposed [xia2024gsva, lai2023lisa, zhang2024psalm, rasheed2024glamm], sparking interest in segmenting and tracking objects in videos, i.e., Reasoning and Referring Video Object Segmentation (RVOS). Although initial architectures for RVOS involved using text and vision encoders coupled with a Transformer decoder to regress masks [yan2023mutr, botach2022mttr, wu2022referformer, bellver2020refvos, seo2020urvos], recent approaches [bai2024videolisa, yan2024visa, yuan2025sa2va] adopt MLLM-based architectures to better reason about complex text prompts and to combine multiple tasks. This is usually done by extending a Multimodal LLM architecture with a downstream segmentation model and by predicting a special segmentation token for each target object. This token is then input into the segmentation model together with video frame features to predict masks. For the video segmentation task, the image-first PSALM model [zhang2024psalm] can be conditioned on a previous mask to segment the next frame. Similarly, TrackGPT [zhu2023trackgpt] processes a single frame at a time. VISA [yan2024visa] predicts a single token for a carefully selected keyframe in a video and tracks that corresponding object. VideoLISA [bai2024videolisa] uses SAM [kirillov2023sam] and reuses a single token for multiple frames. More recently, VideoGLaMM [munasinghe2024videoglamma] and Sa2VA [yuan2025sa2va] use the SAM2 [ravi2025sam2] segmentation model, but do not finetune the memory components. In this work, we identify that Sa2VA operates differently during inference and training, which harms the performance. To solve this, we present Sa2VA-i, which ensures consistent training and inference and considerably outperforms Sa2VA.
@@
Works that use Sa2VA  Sa2VA has rapidly become a foundational model for language-grounded segmentation, and has been widely adopted and extended across diverse research efforts. For instance, its capabilities are leveraged for dataset creation, with AnimeShooter [qiu2025animeshooter] using Sa2VA to extract masks of animated characters. Furthermore, several new benchmarks, including EASG-Bench [rodin2025easgbench], FinChart-Bench [shu2025finchartbench] and V-STaR [cheng2025vstar], evaluate Sa2VA as a state-of-the-art method for spatial grounding. The model is also a baseline for comparison to SpatialReasoner-R1 [shen2025finegrained], Omni-R1 [zhong2025omnir1], SeC [zhang2025sec], and DenseWorld-1M [li2025denseworld1m], EgoMask [liang2025egomask]. InterRVOS [jin2025interrvos], Dense360 [zhou2025dense360], UniBiomed [wu2025unibiomed], DeSa2VA [jisheng2025desa2va], and SAMA [sun2025sama] are built directly on top of Sa2VA. The practical utility of Sa2VA is evident in applications such as Gondola [chen2025gondola] and FineGrasp [du2025finegrasp], which use it to segment objects for grasping. Furthermore, Sa2VA prominence was highlighted during the PVUW MeViS Challenge at CVPR 2025 [ding2024pvuw], where the first [fang2025pvuw1st] and third [yuan2025pvuw3rd] place solutions were based on Sa2VA. The frequent use of Sa2VA in the literature highlights the relevance of this work in which we identify and solve the inference-training inconsistency in Sa2VA. With Sa2VA-i, we provide a stronger base model for various benchmarks, applications and downstream tasks. We encourage authors to apply our suggested improvements to ensure proper comparisons and improve their results.
@@ 3.1 Preliminaries
SAM and SAM2  SAM [kirillov2023sam] is an interactive image segmentation model that consists of image encoder, prompt encoder and mask decoder. Box, point or mask inputs from the user are encoded by the prompt encoder into a single feature vector. Using this feature vector, as well as image features encoded by the image encoder, the mask decoder generates a segmentation output for the selected object. For interactive video segmentation, SAM2 [ravi2025sam2] augments the original SAM architecture with several memory components. SAM2 processes videos in a streaming fashion, i.e., frame by frame. Like SAM, SAM2 segments objects based on a box prompt, mask prompt or point prompt, which are provided for one of the video frames in this case. Given an input frame, SAM2 first generates per-frame features with its image encoder. Then, SAM2 applies memory attention to condition these per-frame features on the features from earlier frames in which an object was segmented, which were stored in a memory bank. Subsequently, the mask decoder predicts a segmentation mask based on the conditioned frame and the input prompts. Finally, the image and object features are encoded using a memory encoder and stored in the memory bank such that they can be used to condition future frames.
@@
For referring video segmentation, Sa2VA also predicts only a single [SEG] token, and reuses this token to predict masks for multiple video frames, similar to VideoLISA [bai2024videolisa]. Due to the limited context window of the MLLM, Sa2VA samples $T$ video frames, only feeds the visual tokens for those $T$ frames through the MLLM, and makes segmentation predictions with SAM2 only for those frames. To obtain predictions for all $I>T$ frames in a video, Sa2VA applies post-processing using the same SAM2 decoder.
@@
Method MLLM Frames MeViS [ding2023mevis] Ref-YT-VOS [seo2020urvos] Ref-DAVIS17 [khoreva2018rdavis]
@@
LISA [lai2023lisa] LLaVA-7B 37.2 35.1 39.4 53.9 53.4 54.3 64.8 62.2 67.3
LISA [lai2023lisa] LLaVA-13B 37.9 35.8 40.0 54.4 54.0 54.8 66.0 63.2 68.8
TrackGPT [zhu2023trackgpt] LLaVA-7B 40.1 37.6 42.6 56.4 55.3 57.4 63.2 59.4 67.0
TrackGPT [zhu2023trackgpt] LLaVA-13B 41.2 39.2 43.1 59.5 58.1 60.8 66.5 62.7 70.4
PSALM [zhang2024psalm] Phi-1.5 68.8 65.9 71.7
VISA [yan2024visa] Chat-UniVi-7B 43.5 40.7 46.3 61.5 59.8 63.2 69.4 66.3 72.5
VISA [yan2024visa] Chat-UniVi-13B 44.5 41.8 47.1 63.0 61.4 64.7 70.4 67.0 73.8
VideoLISA [bai2024videolisa] LLaVa-Phi-3-V-3.8B 44.4 41.3 47.6 63.7 61.7 65.7 68.8 64.9 72.7
InstructSeg [wei2024instructseg] Mipha-3B 67.5 65.4 69.5 71.1 67.3 74.9
VideoGLaMM [munasinghe2024videoglamma] Phi3-Mini-3.8B 45.2 42.0 48.2
SAMWISE [cuttano2024samwisea] RoBERTa 49.5 46.6 52.4 69.2 67.8 70.6 70.6 67.4 74.5
VRS-HQ [gong2025vrshq] Chat-UniVi-13B 50.9 48.0 53.7 71.0 69.0 73.1 74.4 71.0 77.9
GLUS [lin2025glus] LLaVA-7B 51.3 48.5 54.2 67.3 65.5 69.0
MPG-SAM 2 [rong2025mpgsam] BEiT 53.7 50.7 56.7 73.9 71.7 76.1 72.4 68.8 76.0
Sa2VA [yuan2025sa2va] InternVL2.5-1B 5 47.0 68.0 69.5
Sa2VA [yuan2025sa2va] InternVL2.5-4B 5 46.4 43.3 49.5 71.3 69.1 73.5 73.7 69.6 77.8
Sa2VA [yuan2025sa2va] InternVL2.5-8B 5 51.5 72.3 75.9
Sa2VA [yuan2025sa2va] InternVL2.5-26B 5 52.1 75.1 78.6
@@
Table 2: Reasoning Video Segmentation on ReVOS [yan2024visa]. Sa2VA-i outperforms Sa2VA and other methods by a large margin, obtaining state-of-the-art performance.
@@
VISA [yan2024visa] Chat-UniVi-7B 39.2 36.7 41.7 52.9 51.1 54.7 46.1 43.9 48.2
VISA [yan2024visa] Chat-UniVi-13B 40.9 38.3 43.5 54.1 52.3 55.8 47.5 45.3 49.7
InstructSeg [wei2024instructseg] Mipha-3B 51.9 49.2 54.7 57.0 54.8 59.2 54.5 52.0 56.9
VRS-HQ [gong2025vrshq] Chat-UniVi-13B 56.8 54.1 59.4 63.3 61.1 65.5 60.0 57.6 62.5
Sa2VA [yuan2025sa2va] InternVL2.5-4B 56.2 53.2 59.1 63.0 60.5 65.5 59.6 56.8 62.3
@@ 4.1 Experimental Setup
Evaluation  We evaluate Sa2VA-i using multiple popular benchmarks for video-level referring and reasoning segmentation [yan2024visa, khoreva2018rdavis, ding2023mevis, seo2020urvos]. We use the standard $\mathcal{J}\&\mathcal{F}$ metric, which is the geometric mean of the Jaccard Score, $\mathcal{J}$, and the Boundary F-Score, $\mathcal{F}$ [perazzi2016davis]. We use the MeViS dataset [ding2023mevis] for ablations, and report scores obtained from the public evaluation server on the val split.
@@ 4.2 Comparison to State of the Art
We compare Sa2VA-i with existing state-of-the-art referring video segmentation methods on multiple benchmarks. As shown in Tab. 1 and Tab. 2, Sa2VA-i achieves state-of-the-art results on all reported benchmarks and improves upon Sa2VA results using the same checkpoints and the same number of frames. Sa2VA-i’s performance improvement is particularly substantial on the challenging MeVIS benchmark [ding2023mevis]. It scores +10.2 $\mathcal{J}$$\mathcal{F}$ higher for the 4B model, +8.0 $\mathcal{J}$$\mathcal{F}$ for the 8B model, and +11.1 $\mathcal{J}$$\mathcal{F}$ for the 26B model. We attribute performance improvements to the improved consistency between training and inference, as well as better frame sampling. Additionally, we observe that the number of sampled frames during testing has a large influence on the results, but that the optimal value depends on the dataset.
@@ 4.3 Experiments
Table 3: Consistent Training and Inference. Using Sa2VA’s original model weights, we show gradual improvements that come from using our improved Sa2VA-i version with consistent training and inference. We further improve performance by using uniform sampling during inference instead of sampling first frames. Evaluation on MeViS val [ding2023mevis].
@@
Table 4: Frame Sampling. We evaluate the impact of different frame sampling strategies for both Sa2VA and Sa2VA-i. Uniform sampling consistently improves performance, with the largest gains observed when training and inference sampling are matched. Evaluation on MeViS val [ding2023mevis].
@@
Frame Sampling  We experiment with different frame sampling strategies during both training and inference, and report the results in Tab. 4. Similar to other works [fang2025pvuw1st, yuan2025pvuw3rd], we observe improved performance when using uniform sampling during inference. To ensure training-inference consistency and to see the possible gains from training the model with uniform sampling, we train a 4B-parameter model with 8 frames sampled randomly and uniformly. During training with uniform sampling, we apply a random offset from the start to increase variation in the frames seen during training. We observe that uniform sampling and consistent sampling strategy yields even better performance, even for the original Sa2VA model.
Table 5: Number of Sampled Frames. Using Sa2VA’s original model weights, we show that gradual improvements can be obtained by using more frames during inference. Performance is progressively improved, with diminishing returns when using more then 12 frames. Evaluation on MeViS val [ding2023mevis].
@@
Number of Sampled Frames  Tab. 5 evaluates the impact of increasing the number of sampled frames that are fed to the MLLM and used to make the initial segmentation predictions, with the same model that has seen 5 frames during training. We find that increasing the number of sampled frames improves performance by non-negligible margins. In particular, the MeVIS performance can be improved by +1.2 $\mathcal{J}$$\mathcal{F}$ when using 12 sampled frames instead of 5. These findings are similar to the results reported in [fang2025pvuw1st], where Sa2VA obtains higher scores when using more frames. However, while the original Sa2VA appears to benefit from using more frames without performance degradation, we observe diminishing returns for Sa2VA-i after sampling more than 12 frames.
@@
Table 6: Overprompting on Ref-DAVIS17 [khoreva2018rdavis]. We vary the number of frames processed by the MLLM and the number of initially predicted masks used to prompt SAM2 for mask propagation. On this dataset, where the target is typically visible in the first frame, performance improves as the MLLM observes more frames, while prompting SAM2 with multiple initial masks degrades performance.
@@
The 7th LSVOS challenge at ICCV 2025 features two distinct tracks: a Video Object Segmentation (VOS) track evaluated on the complex LVOS and MOSE datasets [MOSE, MOSEv2], and a Referring Video Object Segmentation (RVOS) track utilizing the motion-focused MeViS benchmark [ding2023mevis]. As shown in Tab. 7, our method achieved third place in the competition on the RVOS track, demonstrating strong performance in addressing these challenging scenarios. We achieve this result by using 12 frames during inference.